OpenStudy (anonymous):

Solve the following inequalities, give the solution in set notation, and describe (using words) the solution graph. N - 5 1/2>3/4(N-6)

OpenStudy (anonymous):

\[N-5\frac{ 1 }{ 2 }\ > \frac{ 3 }{ 4 }\left( N-6 \right)\] \[N-\frac{ 11 }{ 2 }>\frac{ 3 }{ 4 }\left( N-6 \right)\] 4N-11>3(N-6) complete it.